# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v1

@@ B.3 Computation Cost
Table 10 summarizes the per-task trainable parameters of various continual-learning methods. Standard LoRA [32], CoDyRA [48], and O-LoRA [74] each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix. Mixture-of-Experts variants such as MoE-LoRA and MoE-Adapter [85] additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts. LB-CL [58] introduce $r$ additional parameters, mimic the singular values of SVD.
@@
By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA. Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one-third of the activated parameters of a standard LoRA.
@@
B.4 Extended Visualizations of Rank Activations
@@
MoRA retains task-specific semantics without forgetting. Fig. 5 shows the activation maps for the same Task 1 image after training on Task 1 (a), Task 2 (b), and the last Task 10 (c). Patches corresponding to the airplane object are outlined in orange. In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane-specific knowledge into Rank 0. Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations. In other words, MoRA effectively memorizes and preserves task-relevant semantics, thereby mitigating catastrophic forgetting.
MoRA encodes generic semantics that are reused across tasks. Fig. 6 examines an input image from Task 9 before and after learning Task 9. Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky-background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank. In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data. Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics. This persistent reuse of Rank 11 across tasks illustrates MoRA’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
@@
B.5 Statistical Analyses of Contributing Rank Activations
@@
Broader Impacts
MoRA’s self-activated, sparse mixture-of-ranks method makes continual adaptation of large vision–language and language models far more efficient and reliable. By updating only the most relevant low-rank subspaces, it reduce compute and memory requirements, enabling on-device personalization in domains like education, healthcare, and finance. This democratizes access to powerful, continually evolved models for smaller teams and resource-constrained settings, driving innovation and improving quality of life. Preserving pre-trained knowledge also ensures stability in safety-critical systems, such as medical diagnostics and autonomous vehicles, where forgetting of pre-trained capabilities can be dangerous.
The focus of this paper is fundamental research, and is broadly applicable to model fine-tuning techniques. Thus, it is possible that MoRA’s lightweight updates could be misused to insert stealthy backdoors or reinforce hidden biases that are hard to detect. Because it leaves most pre-trained parameters untouched, existing biases may persist or become entrenched. To mitigate these risks, we recommend strict audit logging of rank-level updates, anomaly detection on sparse changes, robust access controls, and routine bias and fairness assessments alongside any deployment.

### v2

@@ A.7 Computation Cost
Table 10 summarizes the per‐task trainable parameters of various continual‐learning methods. Standard LoRA (Hu et al., 2021), CoDyRA (Lu et al., 2024), and O-LoRA (Wang et al., 2023) each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix. Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter (Yu et al., 2024) additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts. LB-CL (Qiao & Mahdavi, 2024) introduce $r$ additional parameters, mimic the singular values of SVD.
@@
By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA. Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
Trainable parameters and training GPU memory. Beyond the estimated parameter counts in Table 10, in Table 15, we measured the actual trainable parameters for each continual‐learning task and GPU memory usage, under the same settings as Table 1 in the main paper.
LWF (Li & Hoiem, 2017) and ZSCL (Zheng et al., 2023) perform full‐parameter fine‐tuning, consuming the most parameters and memory. MoE‐Adapters (Yu et al., 2024) maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor. CoDyRA (Lu et al., 2024) trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters. MoRA introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low gpu memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
@@
A.8 Extended Visualizations of Rank Activations
@@
MoRA retains task‐specific semantics without forgetting. Fig. 4 shows the activation maps for the same Task 1 image after training on Task 1 (a), Task 2 (b), and the last Task 10 (c). Patches corresponding to the airplane object are outlined in orange. In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane‐specific knowledge into Rank 0. Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations. In other words, MoRA effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
MoRA encodes generic semantics that are reused across tasks. Fig. 5 examines an input image from Task 9 before and after learning Task 9. Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank. In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data. Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics. This persistent reuse of Rank 11 across tasks illustrates MoRA’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
@@
A.9 Statistical Analyses of Contributing Rank Activations